# SD2: Self-Distilled Sparse Drafters

### v1

@@ Abstract
Speculative decoding is a powerful technique for reducing the latency of Large Language Models (LLMs), offering a fault-tolerant framework that enables the use of highly compressed draft models. In this work, we introduce Self-Distilled Sparse Drafters (SD2), a novel methodology that leverages self-data distillation and fine-grained weight sparsity to produce highly efficient and well-aligned draft models. SD2 systematically enhances draft token acceptance rates while significantly reducing Multiply-Accumulate operations (MACs), even in the Universal Assisted Generation (UAG) setting, where draft and target models originate from different model families. On a Llama-3.1-70B target model, SD2 provides a $\times$1.59 higher Mean Accepted Length (MAL) compared to layer-pruned draft models and reduces MACs by over 43.87% with a 8.36% reduction in MAL compared to a dense draft models. Our results highlight the potential of sparsity-aware fine-tuning and compression strategies to improve LLM inference efficiency while maintaining alignment with target models.
@@ 1 Introduction
Large Language Models (LLMs) have proven to have high utility in a wide variety of contexts. However, the causal dependency between preceding and subsequent tokens results in approximately ×$\times 10$ higher latency for sequence generation compared to processing an equivalent length sequence in parallel (Liu et al., 2023). The high computational cost of LLMs has motivated significant research into methods which improve their efficiency, including: quantization (Gholami et al., 2021; Kurtic et al., 2024), pruning (Ma et al., 2023; Gromov et al., 2024), weight sparsity (Frantar and Alistarh, 2023; Sun et al., 2023; Yin et al., 2023b), activation sparsity (Mirzadeh et al., 2023), KV-cache compression (Zhang et al., 2024b), distillation (Kim and Rush, 2016; Hsieh et al., 2023), and matrix decomposition (Hu et al., 2021; Liu et al., 2024b). However, these methods typically trade improved efficiency for decreased model quality. (Yin et al., 2023a; Jaiswal et al., 2023).
@@
Figure 2: SpecBench MAL for SD2 Qwen-2.5 models drafting for Llama-3.1-70B-Instruct in the UAG setting. These results illustrate the benefits of SD2 for aligning draft models even across different model families. SD2 Qwen drafters achieve a higher MAL than their dense counterparts.
@@
Speculative decoding is a uniquely well-suited setting for using sparse draft models since any errors produced during drafting can be gracefully recovered from during the verification step. Other compression strategies such as quantization succeed in reducing memory overhead, but may not reduce latency (Kurtic et al., 2024). Performing operations with parameters quantized to low bit-width data types requires dequantization to a hardware-native data type, typically 8- or 16-bit types for presently available accelerators. As a result, quantization may lead to increased latency in many circumstances, particularly for relatively small draft models which may not be memory-bound. In contrast, SNNs reduce the total amount of Multiply-Accumulate operations (MACs) required per forward pass which, in an ideal setting, would correspond directly with latency improvements. In practice, accessing non-contiguous sparse parameters and storing the associated sparse structure metadata can lead to significant overhead on Graphics Processing Units (GPUs) (Hooker, 2020). Achieving real-world acceleration for fine-grained SNNs requires specialized kernels (Neural Magic, 2021; Schultheis and Babbar, 2023; Lasby et al., 2023; Frantar et al., 2024) and hardware such as Cerebras’ wafer-scale engine. Thangarasa et al. (2023b); Hall et al. (2023); Lie (2023); Agarwalla et al. (2024).
Given the practical limitations of fine-grained sparsity, structured pruning — the removal of layers, neurons, or other substructures — of LLMs to obtain performant draft models is a more hardware-friendly alternative. In particular, pruning entire transformer blocks from the decoder of LLMs provides low latency models which retain their quality under moderate compression ratios (Liu et al., 2023; Gromov et al., 2024; Men et al., 2024; Kim et al., 2024; Sun et al., 2024b). A natural extension of these works is to leverage layer-pruning techniques to obtain draft models as examined by  Thangarasa et al. (2024). However, it is not clear a priori if the model quality degradation of layer-pruned models can be overcome by their improved latency for use as draft models in the speculative decoding setting.
In this work, we propose Self-Distilled Sparse Drafters (SD2)— a novel methodology for obtaining efficient, well-aligned draft models — by utilizing self-data distillation fine-tuning (Yang et al., 2024b; Thangarasa et al., 2024) and weight sparsity. Specifically, we make the following contributions:
@@
  • We showcase the effectiveness of self-data distillation fine-tuning for model alignment, even when aligning with a different model family in the Universal Assisted Generation (UAG) setting;
  • We find that the theoretical end-to-end acceleration of speculative decoding when using fine-grained sparse draft models exceeds that of dense or layer-pruned draft models, but existing optimized sparse representations fail to offer practical acceleration.
@@ 2.1 Speculative decoding
$\tilde{x}_{n+j}=\operatorname*{arg\,max}\;p_{j}\quad\forall{j}\in\{1,...,k\}$ (1)
$x_{n+j}=\begin{cases}\tilde{x}_{n+j}&\text{if $\tilde{x}_{n+j}=\operatorname*{arg\,max}\;q_{n+j}$}\\ \operatorname*{arg\,max}\;q_{n+j}&\text{otherwise}\end{cases}$ (2)
@@
Improvement $\text{Improvement Factor}=\frac{\text{MAL}}{kc+1}$ (3)
@@ 2.2 Self-data distillation
Self-data distillation consists of generating a synthetic dataset, $D_{self}$, whose outputs are generated by a model of interest. We synthetically curate fine-tuning datasets following  Yang et al. (2024c); Thangarasa et al. (2024). The distilled labels are generated by the target model, $M_{t}$ based on the input sequences, ground truth labels, and task-specific context of one or more supervised fine-tuning datasets $D_{f}$. Specifically, given a task specific context $\mathbf{C^{t}}$, original input sequence $\mathbf{X^{t}}$, and original ground truth label $\mathbf{Y^{t}}$. We combine these components into a new input sequence $\mathbf{X^{\prime}}$.
$\mathbf{\tilde{Y}}=M_{t}(\mathbf{X^{\prime}})\quad\text{where}\;\mathbf{X}^{\prime}=\mathbf{C^{t}}\mathbin{\|}\mathbf{X^{t}}\mathbin{\|}\mathbf{Y^{t}}$ (4)
In the original formulation of self-data distillation, Yang et al. (2024c) only extract the distilled label $\mathbf{\tilde{Y}}$ if and only if it aligns with the original ground truth label, ${\mathbf{Y^{t}}}$. For instance, only accepting the distilled label if it is mathematically equivalent to the ground truth label. However, for tasks that do not yield a closed-form solution, it can be challenging to assess the validity of the distilled label.
@@ 2.3 Fine-grained pruning
To obtain our fine-grained sparse draft models, $M_{d}$, we use SparseGPT (Frantar and Alistarh, 2023). However, we emphasize that our method is compatible with any pruning algorithm. In addition to the original SparseGPT hyperparameter settings, we experiment with non-uniform layer-wise sparsity distributions such as Outlier Weighted Layer-wise sparsity (OWL) (Yin et al., 2023b) and a novel distribution inspired by the angular cosine distance measure from  Gromov et al. (2024). See Appendix B for a discussion of our findings. Ultimately, we use a uniform layer-wise sparsity distribution for all results in this paper.
@@ 2.4 Sparse fine-tuning
In typical supervised fine-tuning settings, a dense LLM $M$ with parameters $\theta$ is fine-tuned using a supervised dataset $D_{f}$ containing input sequences $\mathcal{X}$ and ground truth output sequences $\mathcal{Y}$. For each token, $y_{j}$, in the ground truth output sequences, $\mathbf{Y_{i}}\in\mathcal{Y}$, the model outputs a probability distribution over its vocabulary conditioned by the input sequences, $\mathbf{X_{i}}\in\mathcal{X}$, any output tokens preceding the current token position, and the model parameters: $Q({y}_{j}|\mathbf{X_{i}},\{y_{1},...,y_{j-1}\},\theta)$ where $Q(\cdot)$ represents the raw model logits normalized with the softmax function. The total loss per mini-batch is the average negative log-likelihood across all sequences and tokens:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\frac{1}{S_{i}}\sum_{j=1}^{S_{i}-1}P(y_{j})\log(Q({y}_{j}|\mathbf{X_{i}},y_{\leq j-1},\theta))$ (5)
@@ 2.5 Layer pruning
$d(x_{D_{t}}^{i}\text{,}\;x_{D_{t}}^{i+n})=\frac{1}{\pi}\text{arccos}\left(\frac{x_{D_{t}}^{i}\cdot x_{D_{t}}^{i+n}}{\left\|x_{D_{t}}^{i}\right\|\left\|x_{D_{t}}^{i+n}\right\|}\right)$ (6)
@@ 3 Results
To empirically evaluate SD2 and our primary hypothesis that sparse draft models can outperform dense or layer-pruned models, we evaluate a variety of draft model candidates across the Llama-3.2 (Llama Team, AI @ Meta, 2024) and Qwen-2.5 (Qwen et al., 2025) model families. For Llama, we evaluate drafting with Llama-3.2 1B-Instruct and 3B-Instruct for Llama-3.1-70B-Instruct. For Qwen-2.5, we evaluate 0.5B-Instruct, 1.5B-Instruct, 3B-Instruct drafting for Qwen-2.5-72B-Instruct. We also consider the UAG setting in which the Qwen draft models are aligned to and evaluated with a Llama-3.1-70B-Instruct target model.
We pruned and fine-tuned fine-grained sparse draft models with both unstructured and 2:4 sparsity. We use a uniform layer-wise sparsity distribution and prune all decoder blocks, excluding the embedding and lm-head layers. The sparsity levels reported in our results refer to the overall sparsity of the decoder. As baselines, we compare these candidates with 50% layer-pruned [^2] Specifically, 50% of the decoder blocks are pruned. and dense models. See Appendix A for hyperparameters and implementation details.
@@ 3.1.1 OpenLLM leaderboard V1 benchmarks
We evaluate our models using the default multi-shot settings on the OpenLLM Leaderboard V1 tasks using the EleutherAI evaluation harness (Gao et al., 2023). These tasks include: 25-shot ARC-C (Clark et al., 2018), 5-shot strict exact match GSM8k (Cobbe et al., 2021), 10-shot HellaSwag (Zellers et al., 2019), 5-shot MMLU (Hendrycks et al., 2021), 5-shot Winogrande (Sakaguchi et al., 2019), and 0-shot multi-true (MC2) TruthfulQA (Lin et al., 2022). We report byte-length normalized accuracies for ARC-C and HellaSwag [^3] Reported as the acc_norm field in the EleutherAI evaluation harness outputs. See  Gao (2021) for more details..
See Figure 3 for the mean accuracy for models pruned and fine-tuned using the SD2 methodology. We find that one-shot pruned models suffer high degradation on the tasks evaluated; however, after fine-tuning the fine-grained sparse models recover much of their accuracy, particularly the 50% unstructured and 2:4 sparse models. For the Qwen-2.5 1.5 and 3B models, we find that the unstructured 50% models approach or exceed the accuracy of the dense baseline. Notably, the 50% layer-pruned models suffer high degradation, even after fine-tuning. See Appendix E for more detailed results.
@@
Figure 4: MAL for Llama-3.2 and Qwen-2.5 model families on SpecBench (Xia et al., 2024) for layer-pruned, unstructured sparse, and 2:4 sparse draft models. Lighter and darker shades show results after one-shot pruning and SD2 fine-tuning, respectively. Dense model baseline are depicted in horizontal grey lines on each plot. Top row: Llama-3.2 draft models speculating for a Llama-3.1-70B-Instruct target model. Bottom row: Qwen-2.5 draft models speculating for Qwen-2.5-72B-Instruct. Across both model families, we observe that the fine-grained sparse draft models achieve significantly higher MALs than the layer-pruned models.
@@ 3.1.2 Speculative decoding
SpecBench Xia et al. (2024) assesses draft token Acceptance Rate (AR) across a variety of tasks including translation, summarization, question answering, mathematical reasoning, multi-turn conversation, and retrieval augmented generation (RAG).
@@
The importance of the draft model alignment is clear; our self-data distillation and fine-tuning process enables much higher MAL than the one-shot pruned models. Even more strikingly, the SD2 effectively aligns the draft model in the UAG setting across model families, as can be seen in Figure 2. Remarkably, the sparse Qwen drafters outperform their dense counterparts under this setting, further highlighting the importance of draft model alignment and the benefits of SD2.

### v2

@@ Abstract
Speculative decoding is a powerful technique for reducing the latency of Large Language Models (LLMs), offering a fault-tolerant framework that enables the use of highly compressed draft models. In this work, we introduce Self-Distilled Sparse Drafters (SD2), a novel methodology that leverages self-data distillation and fine-grained weight sparsity to produce highly efficient and well-aligned draft models. SD2 systematically enhances draft token acceptance rates while significantly reducing Multiply-Accumulate operations (MACs), even in the Universal Assisted Generation (UAG) setting, where draft and target models originate from different model families. On a Llama-3.1-70B target model, SD2 provides a 1.59$\times$ higher Mean Accepted Length (MAL) compared to layer-pruned draft models and reduces MACs by over 43.87% with a 8.36% reduction in MAL compared to a dense draft models. Our 1.5B and 3B unstructured sparse drafters outperform both dense and layer-pruned models in terms of end-to-end latency improvements; highlighting the potential of sparsity-aware fine-tuning and compression strategies to improve LLM inference efficiency while maintaining alignment with target models.
@@ 1 Introduction
Large Language Models (LLMs) have proven to have high utility in a wide variety of contexts. However, the causal dependency between preceding and subsequent tokens results in approximately 10$\times$ higher latency for sequence generation compared to processing an equivalent length sequence in parallel (Liu et al., 2023). The high computational cost of LLMs has motivated significant research into methods which improve their efficiency, including: quantization (Gholami et al., 2022; Kurtic et al., 2024), pruning (Ma et al., 2023; Gromov et al., 2025), weight sparsity (Frantar and Alistarh, 2023; Sun et al., 2024b; Yin et al., 2024b), activation sparsity (Mirzadeh et al., 2023), KV-cache compression (Zhang et al., 2024c), distillation (Kim and Rush, 2016; Hsieh et al., 2023), and matrix decomposition (Hu et al., 2021; Liu et al., 2024b). However, these methods typically trade improved efficiency for decreased model quality. (Yin et al., 2024a; Jaiswal et al., 2024).
@@
Figure 1: Improvement factor of dense, layer-pruned, and SD2 unstructured Llama and Qwen models drafting for Llama-3.1-70B-Instruct and Qwen-2.5-72B-Instruct, respectively. SD2 drafters outperform layer-pruned draft models and dense drafters in the 1.5B and 3B model size categories.
(a) UAG Qwen drafting for Llama
(b) MACs analysis
Figure 2: (2(a)): SpecBench MAL for SD2 Qwen-2.5 models drafting for Llama-3.1-70B-Instruct in the UAG setting. These results illustrate the benefits of SD2 for aligning draft models even across different model families. SD2 Qwen drafters achieve a higher MAL than their dense counterparts. (2(b)): MAL vs. MACs for layer-pruned and sparse draft models. Particularly notable are the Qwen-2.5 unstructured sparse drafters which approach iso-MAC performance compared to the dense models.
@@
Speculative decoding is a uniquely well-suited setting for using sparse draft models since any errors produced during drafting can be gracefully recovered from during the verification step. Other compression strategies such as quantization succeed in reducing memory overhead, but may not reduce latency (Kurtic et al., 2024). Performing operations with parameters quantized to low bit-width data types requires dequantization to a hardware-native data type, typically 8- or 16-bit types for presently available accelerators. As a result, quantization may lead to increased latency in many circumstances, particularly for relatively small draft models which may not be memory-bound. In contrast, SNNs reduce the total amount of Multiply-Accumulate operations (MACs) required per forward pass which, in an ideal setting, would correspond directly with latency improvements. In practice, accessing non-contiguous sparse parameters and storing the associated sparse structure metadata can lead to significant overhead on Graphics Processing Units (GPUs) (Hooker, 2020). Achieving real-world acceleration for fine-grained SNNs requires specialized kernels (Neural Magic, 2021; Schultheis and Babbar, 2023; Lasby et al., 2023; Frantar et al., 2025) and hardware such as Cerebras’ wafer-scale engine (Thangarasa et al., 2024; Hall et al., 2023; Lie, 2023; Agarwalla et al., 2024).
Given the practical limitations of fine-grained sparsity, structured pruning — the removal of layers, neurons, or other substructures — of LLMs to obtain performant draft models is a more hardware-friendly alternative. In particular, pruning entire transformer blocks from the decoder of LLMs provides low latency models which retain their quality under moderate compression ratios (Liu et al., 2023; Gromov et al., 2025; Men et al., 2024; Kim et al., 2024; Sun et al., 2025). A natural extension of these works is to leverage layer-pruning techniques to obtain draft models as examined by  Thangarasa et al. (2025). However, it is not clear a priori if the model quality degradation of layer-pruned models can be overcome by their improved latency for use as draft models in the speculative decoding setting.
In this work, we propose Self-Distilled Sparse Drafters (SD2)— a novel methodology for obtaining efficient, well-aligned draft models — by utilizing self-data distillation fine-tuning (Yang et al., 2024b; Thangarasa et al., 2025) and weight sparsity. Specifically, we make the following contributions:
@@
  • We showcase the effectiveness of self-data distillation fine-tuning for model alignment, even when aligning with a different model family with Universal Assisted Generation (UAG);
  • When paired with optimized sparse representations, we find that the end-to-end acceleration of speculative decoding with fine-grained sparse draft models is comparable to and in some cases exceeds that of dense draft models, particularly with larger draft model sizes.
@@ 2.1 Speculative decoding
$\tilde{x}_{n+j}=\operatorname*{arg\,max}\;p_{j}\quad\forall{j}\in\{1,...,k\},$ (1)
$x_{n+j}=\begin{cases}\tilde{x}_{n+j}&\text{if $\tilde{x}_{n+j}=\operatorname*{arg\,max}\;q_{n+j}$}\\ \operatorname*{arg\,max}\;q_{n+j}&\text{otherwise}.\end{cases}$ (2)
@@
Improvement $\text{Improvement Factor}=\frac{\text{MAL}}{kc+1},$ (3)
@@ 2.2 Self-data distillation
Self-data distillation consists of generating a dataset, $D_{self}$, whose outputs are generated by a model of interest. We curate fine-tuning datasets following  Yang et al. (2024b); Thangarasa et al. (2025). The distilled labels are generated by the target model, $M_{t}$ based on the input sequences, ground truth labels, and task-specific context of one or more supervised fine-tuning datasets $D_{f}$. Specifically, given a task specific context $\mathbf{C^{t}}$, original input sequence $\mathbf{X^{t}}$, and original ground truth label $\mathbf{Y^{t}}$. We combine these components into a new input sequence $\mathbf{X^{\prime}}$
$\mathbf{\tilde{Y}}=M_{t}(\mathbf{X^{\prime}})\quad\text{where}\;\mathbf{X}^{\prime}=\mathbf{C^{t}}\mathbin{\|}\mathbf{X^{t}}\mathbin{\|}\mathbf{Y^{t}}.$ (4)
In the original formulation of self-data distillation, Yang et al. (2024b) only extract the distilled label $\mathbf{\tilde{Y}}$ if and only if it aligns with the original ground truth label, ${\mathbf{Y^{t}}}$. For instance, only accepting the distilled label if it is mathematically equivalent to the ground truth label. However, for tasks that do not yield a closed-form solution, it can be challenging to assess the validity of the distilled label.
@@ 2.3 Fine-grained pruning
To obtain our fine-grained sparse draft models, $M_{d}$, we use SparseGPT (Frantar and Alistarh, 2023). However, we emphasize that our method is compatible with any pruning algorithm. In addition to the original SparseGPT hyperparameter settings, we experiment with non-uniform layer-wise sparsity distributions such as Outlier Weighted Layer-wise sparsity (OWL) (Yin et al., 2024b) and a novel distribution inspired by the angular cosine distance measure from  Gromov et al. (2025). While our experiments found that non-uniform layer-wise sparsity improves perplexity, little benefit was observed on downstream evaluations or speculative decoding. See Appendix D for a more thorough discussion of our findings. Ultimately, we use a uniform layer-wise sparsity distribution for all results presented below.
@@ 2.4 Sparse fine-tuning
In typical supervised fine-tuning settings, a dense LLM $M$ with parameters $\theta$ is fine-tuned using a supervised dataset $D_{f}$ containing input sequences $\mathcal{X}$ and ground truth output sequences $\mathcal{Y}$. For each token, $y_{j}$, in the ground truth output sequences, $\mathbf{Y_{i}}\in\mathcal{Y}$, the model outputs a probability distribution over its vocabulary conditioned by the input sequences, $\mathbf{X_{i}}\in\mathcal{X}$, any output tokens preceding the current token position, and the model parameters: $Q({y}_{j}|\mathbf{X_{i}},\{y_{1},...,y_{j-1}\},\theta)$ where $Q(\cdot)$ represents the language modelling head (LM head) logits normalized with the softmax function. The total loss per mini-batch is the average negative log-likelihood across all sequences and tokens:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\frac{1}{S_{i}}\sum_{j=1}^{S_{i}-1}P(y_{j})\log(Q({y}_{j}|\mathbf{X_{i}},y_{\leq j-1},\theta)),$ (5)
@@ 2.5 Layer pruning
$d(x_{D_{t}}^{i}\text{,}\;x_{D_{t}}^{i+n})=\frac{1}{\pi}\text{arccos}\left(\frac{x_{D_{t}}^{i}\cdot x_{D_{t}}^{i+n}}{\left\|x_{D_{t}}^{i}\right\|\left\|x_{D_{t}}^{i+n}\right\|}\right),$ (6)
@@ 3 Results
To empirically evaluate SD2 and our primary hypothesis that sparse draft models can outperform dense or layer-pruned models, we evaluate a variety of draft model candidates across the Llama-3.2 (Llama Team, AI @ Meta, 2024) and Qwen-2.5 (Yang et al., 2025) model families. For Llama, we evaluate drafting with Llama-3.2 1B-Instruct and 3B-Instruct for Llama-3.1-70B-Instruct. For Qwen-2.5, we evaluate 0.5B-Instruct, 1.5B-Instruct, 3B-Instruct drafting for Qwen-2.5-72B-Instruct. We also consider the UAG setting in which the Qwen draft models are aligned to and evaluated with a Llama-3.1-70B-Instruct target model. We prune and fine-tune fine-grained sparse draft models with both unstructured and 2:4 sparsity using a uniform layer-wise sparsity distribution and prune all decoder blocks, excluding the embedding and LM head. The sparsity levels reported in our results refer to the overall sparsity of the decoder. As baselines, we compare these candidates with 50% layer-pruned [^2] Specifically, 50% of the decoder blocks are pruned. and dense models. See Appendix C for hyperparameters and implementation details.
@@ 3.1.1 OpenLLM leaderboard V1 benchmarks
We evaluate our models using the default multi-shot settings on the OpenLLM Leaderboard V1 tasks using the EleutherAI evaluation harness (Gao et al., 2023). These tasks include: 25-shot ARC-C (Clark et al., 2018), 5-shot strict exact match GSM8k (Cobbe et al., 2021), 10-shot HellaSwag (Zellers et al., 2019), 5-shot MMLU (Hendrycks et al., 2021), 5-shot Winogrande (Sakaguchi et al., 2021), and 0-shot multi-true (MC2) TruthfulQA (Lin et al., 2022). We report byte-length normalized accuracies for ARC-C and HellaSwag [^3] Reported as the acc_norm field in the EleutherAI evaluation harness outputs. See  Gao (2021) for details..
See Figure 3 for the mean accuracy for models pruned and fine-tuned using the SD2 methodology. We find that one-shot pruned models suffer high degradation on the tasks evaluated; however, after fine-tuning the fine-grained sparse models recover much of their accuracy, particularly the 50% unstructured and 2:4 sparse models. For the Qwen-2.5 1.5 and 3B models, we find that the unstructured 50% models approach or exceed the accuracy of the dense baseline. Notably, the 50% layer-pruned models suffer high degradation, even after fine-tuning. See Appendix G for more detailed results.
@@
Figure 4: MAL for Llama-3.2 and Qwen-2.5 model families on SpecBench (Xia et al., 2024) for layer-pruned, unstructured sparse, and 2:4 sparse draft models including results after one-shot pruning and SD2 fine-tuning. Dense model baseline are depicted in horizontal grey lines on each plot. Top row: Llama-3.2 draft models speculating for a Llama-3.1-70B-Instruct target model. Bottom row: Qwen-2.5 draft models speculating for Qwen-2.5-72B-Instruct. Across both model families, we observe that the fine-grained sparse draft models achieve significantly higher MALs than the layer-pruned models.
@@ 3.1.2 Speculative decoding
SpecBench Xia et al. (2024) assesses draft token acceptance rate across a variety of tasks including translation, summarization, question answering, mathematical reasoning, multi-turn conversation, and retrieval augmented generation (RAG).
@@
The importance of the draft model alignment is clear; our self-data distillation and fine-tuning process enables much higher MAL than the one-shot pruned models. Even more strikingly, the SD2 effectively aligns the draft model in the UAG setting across model families, as can be seen in Figure 2(a). Remarkably, the sparse Qwen drafters outperform their dense counterparts under this setting, further highlighting the importance of draft model alignment and the benefits of SD2.